The random variable x is known to be uniformly distributed between 4.53 and 9.68. Compute the probability that x is exactly 8.

Group of answer choices

0.674

0.563

0

0.146

1.553

0.326
Answer #1

1) Discrete uniform distribution :

Let a= 4.53 and b= 9.68 then

n= b-a+1 = 6.15

P(x=8) = 1/n

= 1/(6.15)

= 0.163

2) Continuous Uniform distribution :

For any continuous distribution, Probability that x is exactly equal to any constant value (here 8) is 0.

Hence probability is 0.